# A Survey of AI Methods for Geometry Preparation and Mesh Generation in Engineering Simulation

### v1

@@ 7 Block Structured Meshing
Many advanced physics solvers—such as those for shock physics or structural dynamics—depend on meshes composed entirely of quadrilaterals (2D) or hexahedra (3D). Solver accuracy and stability can be highly sensitive to mesh structure, making all-hex and block-structured approaches especially valuable [Benzley1995, Schneider2022]. The principle is straightforward: decompose a complex domain into four-sided or six-sided blocks and fill each with a structured grid, often using transfinite interpolation (TFI) [Thompson1998]. A widely adopted variant, the “pave-and-sweep” strategy [StatenCanannOwen1999] [RocaSarrate2010], generates a quadrilateral surface mesh and extrudes it through a third dimension to form hex elements.
@@
Over decades, geometry- and topology-based methods such as mapping, submapping, and pave-and-sweep have provided the foundation of structured meshing. Production tools embody these strategies—Sandia’s Cubit® [Cubit15_2_UserDoc_2016] integrates an interactive environment for decomposition, while industrial packages like Gridgen [Chawner1995] and GridPro [Ferlemann2010] emphasize smoothing and topology-first multiblock generation for CFD. Yet despite their maturity, these methods leave the hardest step—discovering a valid block decomposition—largely in the hands of expert analysts. Progress in 2D has been tangible, but robust, general-purpose automation in 3D has remained elusive.
@@
The state representation captured local geometry around vertices in fixed-length vectors, while the reward function encouraged quad generation with uniform geometric properties (e.g., area, aspect ratio) and lightly penalized unproductive actions. On axis-aligned datasets restricted to $X$- and $Y$-direction cuts, the agent consistently achieved $100\%$ quadrilateral decompositions. Follow-on work expanded the action set with bisections and generalized the approach to arbitrary polygons [RayRL2025]. For orthogonal shapes, the policy learned to favor edge-aligned cuts that produced rectangles; for general polygons, it generated many quadrilaterals but often left residual triangles. These results show that RL can capture effective decomposition strategies but that fully generalized behavior remains out of reach.
@@
The limitations also highlight a key distinction from standard RL benchmarks: the meshing environment is not static but varies dramatically across CAD models. Capturing such variability with fixed-length features is brittle and often leads to noisy training. Promising directions include richer, multi-scale state representations, larger and more diverse training corpora, and the use of imitation learning [Argall2009SurveyLfD] to provide strong initial policies. In this way, 2D decomposition continues to serve not only as a useful capability in its own right but also as a critical stepping stone toward 3D block-structured and all-hex meshing, where the challenges—and the potential impact—are far greater.
@@
An alternative to the approach taken in [diprete2024reinforcement] is to leave the geometry domain unchanged and build a block structure that discretizes it. This can be achieved by considering the work of Naranyan et. al. [Narayanan2024], who present a learning-based framework for improving the quality of unstructured triangular and quadrilateral meshes. Their approach learns to improve mesh quality via self-play RL with no prior heuristics. They use an actor-only version of the Proximal Policy Optimization (PPO) algorithm (see [schulman2017]) without a critic (i.e., value function) network. They demonstrate their agent’s ability to transform unstructured quadrilateral meshes—obtained by splitting triangles into four quads—into structured meshes that correspond to 2D block decomposition. Although the goal of [Narayanan2024] is not block decomposition, the applicability of their approach opens promising perspectives for pursuing this path.
@@
Complementing these learning-based strategies, Paul et al. [PaulMCTS2025] frame 2D decomposition as Monte Carlo Tree Search (MCTS): starting from the CAD model as root, the search expands candidate partition operations while balancing exploitation and exploration. The method often finds near-optimal quadrilateral decompositions, showing that high-quality layouts can emerge without end-to-end training. Extending MCTS to 3D is especially promising—partial block layouts alone could guide experts and scaffold downstream learning. More broadly, search-based methods complement RL when reward shaping and state design are brittle, and they anticipate the demands of 3D all-hex and block-structured meshing: long-horizon reasoning under sweepability, feature-preservation, and quality constraints, developed next.
@@
All-hex meshing of complex CAD models begins with a sound geometric decomposition. Downstream schemes, including unstructured approaches (for example, pave and sweep), structured mappings (for example, map), and block-structured layouts, then generate conforming, high-quality hexahedra. The central difficulty is not meshing a single sweepable or mappable block, but discovering a valid global decomposition and an executable sequence of operations that (i) respect sweepability constraints, (ii) preserve curvature and sharp features, and (iii) maintain element quality. Decades of practice have codified this workflow in tools such as Sandia’s Cubit®, which provide robust predicates and operations, yet the pivotal decomposition step remains largely manual. This gap motivates AI methods that can capture expert strategies, reason over long action horizons, and adapt decisions to complex geometry.
@@
Recent work has begun to push in this direction. Patel et al. [Patel2021AutoHex] introduced Auto-Hex, which integrates geometric reasoning with RL in a closed loop with Cubit®. Starting from a boundary representation (B-Rep) and chordal-axis skeleton, the system identifies features such as T-junctions, sweep directions, and through-holes. An RL policy then proposes webcut actions conditioned on these cues, with feedback from the meshing environment guiding the agent to favor decompositions that yield hex or hex-dominant meshes. In related 2D work discussed earlier, DiPrete et al. [diprete2024reinforcement] used CAD shape testbeds to study RL-driven cut sequencing and reward design, while Zhang et al. [Zhang2025CADRLHex] frame 3D decomposition as a sequential decision problem and employ Soft Actor-Critic (SAC) to generate sweepable sub-volumes, showing promising decompositions on complex solids while highlighting open challenges in reward shaping and mesh quality.
@@
In parallel, supervised learning has been explored as another route to automation. Quadros et al. [Quadros2025Bipartite] reduce a general 3D CAD model (outer B-Rep “skin”) to a 1D bipartite, heterogeneous graph derived from a 2D chordal-axis skeleton, then apply a graph neural network (GNN) to predict decomposition operations. Graph nodes carry geometric attributes (area, length, centroid), topological attributes (valency, interior loops), and meshing cues (principal sweep directions), with two-way mappings between the B-Rep, skeleton, and graph. Because labeled decompositions are costly, training is performed on synthetic CAD with action labels, yielding $98.72\%$ accuracy in predicting operations such as Cubit® webcuts. Current work is extending this approach to targeted CAD classes and coupling it with RL, as in Auto-Hex, to move beyond labeled datasets and learn the sequencing of webcut commands on general 3D geometries.
@@ 8 Volumetric Parameterizations
Volumetric parameterization maps a solid 3D volume onto simple parametric domain (e.g., a cube or polycube). This mapping must be bijective and low-distortion to support applications such as hexahedral meshing, spline modeling and isogeometric analysis (IGA). While surface parameterization is well-established, extending these concepts to volumes introduces significant challenges due to higher dimensionality as well as complex geometry and topology.
Unlike block-structured meshing, which decomposes geometry into discrete blocks for local meshing, volumetric parameterization constructs a continuous global map from domain to volume. This enables two critical tasks: (i) generating high-quality all-hex meshes for simulation, and (ii) building tensor-product spline domains for IGA, where hex control meshes align naturally with spline-based discretizations [cottrell_isogeometric_2009].
Constructing smooth, bijective, and low-distortion volumetric maps remains an open problem. Complex geometries often preclude simple solutions, and existing methods struggle with element inversion, excessive distortion and extraordinary nodes. These issues directly impact mesh validity and the stability of numerical simulations, particularly for spline-based IGA.
Several frameworks target this problem from a meshing perspective. CubeCover [Nieser2011CubeCover] introduces volumetric frame fields and integer-grid maps to produce boundary-aligned parameterizations. Polycube methods deform the input into a polycube and apply regular grid pullbacks to extract hexahedral elements [Gregson2011Polycube, Fang2016ClosedFormPolycube]. Recent efforts improve robustness with intrinsic formulations [Mandad2022IntrinsicPolycubes] and quantization techniques [Brueckler2022Quantization], enhancing alignment and hex quality.
Beyond meshing, parameterization plays a central role in IGA and geometry processing. Early methods convert unstructured meshes into T-spline representations [wenyan2010a, Wenyan2011a]. Harmonic maps solve Laplace equations with Dirichlet boundaries to parameterize genus-0 and higher-genus volumes [zhang_solid_2012, Wenyan2013c, LeiLiu2012a]. Polycube decomposition has become standard for IGA [LZH2014, Liu2015, HZ2016CMAME, HZL2017], while advanced spline bases—e.g., weighted T-splines [LeiLiu2015a], truncated T-splines [ref:wei16a], and hierarchical B-splines [wei17a, wei2018blended]—are now supported in commercial solvers [Lai2016, Lai2017, yu2022hexgen]. Optimization-based methods introduce Jacobian regularization [Ji2022] and multi-patch schemes to ensure inter-domain continuity [Zou2025].
Despite these advances, robust parameterization for general CAD models remains elusive. Most pipelines still require expert input, difficulty in automation, suffer from high computational costs, and lack guarantees of bijectivity or low distortion—limiting mesh quality and analysis reliability.
Recent AI-driven methods offer a promising shift. Deep learning automates segmentation and feature alignment, transforming hours of manual work into seconds. DeepShape [DeepShapeCode] uses voxelized inputs and CNNs to learn mappings from geometry to polycube domains. DL-Polycube [yu2025dlpolycube] and DDPM-Polycube [yu2025DDPM] predict polycube structures using neural templates or diffusion models. Neural integral parameterization methods model the inverse map directly, maintaining bijectivity and handling boundary constraints [Zhan2025]. Reinforcement learning approaches [TONG2023102109] optimize parameterization strategies that preserve quality across complex inputs. Using genetic algorithms, Evocube [Dumery2022] focuses on the labelling problem, which aims to precompute polycube alignment by assigning one of the base axes to each boundary face of the input. When evaluated on a thousand smooth and CAD meshes, Evocube converges on accurate labels for a wide range of shapes.
Though these methods are still evolving, they demonstrate the feasibility of fully automated, and high-quality parameterization. Neural networks now produce bijective maps, generative models handle topological decomposition, and hybrid AI-classical methods combine learning with robust geometric frameworks.
The potential impact is transformative. Automation can drastically reduce modeling time, improve mesh quality, and lower the barrier for spline-based simulation at scale. While block-structured meshing partitions geometry, volumetric parameterization enables continuous mappings essential for high-fidelity simulation. The integration of AI with geometric theory represents the most scalable path forward.
@@ 9 Parallel Mesh Generation
As simulation fidelity and geometric complexity rise, large-scale FEA increasingly hinges on parallel mesh generation to sustain simulations with hundreds of millions of elements. Misconfigured runtimes can waste substantial wall time at this scale, so choosing the right parameters matters as much as the mesher itself. Both data and computation must be distributed over many processors in an HPC environment. Modern parallel meshers are indispensable, but they expose numerous runtime controls such as block size, work-partitioning strategy, NUMA affinity, and communication thresholds, and these choices have first-order effects on scalability and wall time. Selecting an optimal parameter vector $p$ is a nonlinear and architecture-sensitive problem that has traditionally depended on expert heuristics or costly trial and error.
@@
Analytic performance models exist, but they typically rest on simplifying assumptions and generalize poorly across architectures and problem families [barker2005practical]. Empirical tuning is more faithful but costly: each evaluation consumes a full HPC submission, and queue delays often dominate turnaround time. Collecting even a few dozen labeled runs can take weeks in production settings [angelopoulos2018]. These constraints impede routine optimization when rapid results are needed.
@@
Artificial intelligence offers a complementary path. In particular, supervised learning can train surrogate models that map parameter vectors to execution time [gamatie2019, ipek2005]. Training shifts cost up front. Once learned, the surrogate supports fast what-if exploration over unseen configurations without consuming additional HPC cycles, which enables wide synthetic searches of the parameter space. In what follows we focus on two complementary roles for AI: surrogates that predict performance and enable rapid configuration search, and learned controllers that replace fixed heuristics inside the parallel pipeline.
As a case study, [chrisochoides2025] trains compact feedforward surrogates for a general-purpose parallel mesher that predict execution time from a parameter vector $p$ using small real-world datasets. Two platforms were considered: 172 runs on a shared-memory cc-NUMA system [garner2025numa] and 68 runs on a multi-node distributed-memory system [garner2025dist]. A three-layer network per dataset was sufficient. The trained models then scored 100,000100{,}000 synthetic configurations in under $0.5$ s per case, turning weeks of batch exploration into minutes. On held-out tests the models reached MAPE $2.13\%$ on the shared-memory system and $5.68\%$ on the distributed-memory system [tsolakis2024, garner2025dist]. This accuracy was sufficient to select settings that reduced total execution time. Overall, surrogate-guided search cut tuning time-to-solution by more than $90\%$.
@@
Beyond surrogates, recent work uses learning to decide how parallel meshing is organized. For load balancing in production workflows, trained classifiers select domain-decomposition settings and reduce time-to-solution compared to static heuristics, as shown for snappyHexMesh [Gangopadhyay2025ParCFD]. For decomposition, reinforcement learning with graph neural networks learns partitioners that compete with multilevel methods while exposing a trainable interface to work distribution and communication footprint [Gatti2022RLGNN]. For adaptation, adaptive mesh refinement has been posed as a sequential decision problem, and learned policies make locality-aware refinement choices that scale across partitions and generalize across PDEs and mesh sizes [Yang2023RLAMR]. Taken together, the results show that learned controllers can optimize decomposition, load balancing, and adaptive refinement, while surrogates accelerate search over the remaining parameters.

### v2

@@ 8 Structured Meshing
Many advanced physics solvers, such as those for shock physics or structural dynamics, depend on meshes composed entirely of quadrilaterals (2D) or hexahedra (3D). Solver accuracy and stability can be highly sensitive to mesh structure, making all-hex and block-structured approaches especially valuable [Benzley1995, Schneider2022]. The principle is straightforward: decompose a complex domain into four-sided or six-sided blocks and fill each with a structured grid, often using transfinite interpolation (TFI) [Thompson1998]. A widely adopted variant, the “pave-and-sweep” strategy [StatenCanannOwen1999, RocaSarrate2010], generates a quadrilateral surface mesh and extrudes it through a third dimension to form hex elements.
@@
Over decades, geometry- and topology-based methods such as mapping, submapping, and pave-and-sweep have provided the foundation of structured meshing. Production tools embody these strategies—Sandia’s Cubit [CubitUserDoc1706_2025] integrates an interactive environment for decomposition, while industrial packages like Gridgen [Chawner1995] and GridPro [Ferlemann2010] emphasize smoothing and topology-first multiblock generation for CFD. Yet despite their maturity, these methods leave the hardest step—discovering a valid block decomposition—largely in the hands of expert analysts. Progress in 2D has been tangible, but robust, general-purpose automation in 3D has remained elusive.
Structured meshing also includes field-aligned quadrilateral layouts on surfaces, where the goal is to align elements with global shape and anisotropy rather than to define an explicit multiblock parameterization. Frame fields [Ray2006NSymmetry] provide a compact way to encode these preferred directions, guiding quadrangulation toward coherent alignment with curvature and feature lines while concentrating unavoidable topological transitions into a small set of singularities. Recent learning-based work targets this field construction step directly. NeuFrameQ predicts neural frame fields and uses them to drive scalable anisotropic quadrangulation, emphasizing robustness across diverse shapes [Liu2025NeuFrameQ].
@@
The state representation captured local geometry around vertices in fixed-length vectors, while the reward function encouraged quad generation with uniform geometric properties (e.g., area, aspect ratio) and lightly penalized unproductive actions. On axis-aligned datasets restricted to $X$- and $Y$-direction cuts, the agent consistently achieved $100\%$ quadrilateral decompositions. Follow-on work expanded the action set with bisections and generalized the approach to arbitrary polygons [RayRL2025]. For orthogonal shapes, the policy learned to favor edge-aligned cuts that produced rectangles; for general polygons, it generated mostly quadrilaterals but sometimes left residual triangles. Of course, these triangles could be decomposed into three quadrilateral blocks but those would be sub-optimal. These results show that RL can capture effective decomposition strategies, but fully generalized behavior remains out of reach.
@@
The limitations also highlight a key distinction from standard RL benchmarks: the meshing environment is not static but varies dramatically across CAD models. In particular, the number and type of actions to apply are state-dependent and can vary significantly. Capturing such variability with fixed-length features is brittle and often leads to noisy training. Promising directions include richer, multi-scale state representations, larger and more diverse training corpora, and the use of imitation learning [Argall2009SurveyLfD] to provide strong initial policies. In this way, 2D decomposition continues to serve not only as a useful capability in its own right but also as a critical stepping stone toward 3D block-structured and all-hex meshing, where the challenges—and the potential impact—are far greater.
@@
An alternative to the approach taken in [diprete2024reinforcement] is to leave the geometry domain unchanged and build a block structure that discretizes it. This can be achieved by considering the work of Naranayan et. al. [Narayanan2024], who present a learning-based framework for improving the quality of unstructured triangular and quadrilateral meshes. Their approach learns to improve mesh quality via self-play RL with no prior heuristics. They use an actor-only version of the Proximal Policy Optimization (PPO) algorithm (see [schulman2017]) without a critic (i.e., value function) network. They demonstrate their agent’s ability to transform unstructured quadrilateral meshes—obtained by splitting triangles into four quads—into structured meshes that correspond to 2D block decomposition. Although the goal of [Narayanan2024] is not block decomposition, the applicability of their approach opens promising perspectives for pursuing this path.
@@
Complementing these learning-based strategies, Bourmaud et al. [PaulMCTS2025] frame 2D decomposition as Monte Carlo Tree Search (MCTS): starting from the CAD model as root, the search expands candidate partition operations while balancing exploitation and exploration. The method often finds near-optimal quadrilateral decompositions, showing that high-quality layouts can emerge without end-to-end training. Extending MCTS to 3D is especially promising—partial block layouts alone could guide experts and scaffold downstream learning. More broadly, search-based methods complement RL when reward shaping and state design are brittle, and they anticipate the demands of 3D all-hex and block-structured meshing: long-horizon reasoning under sweepability, feature-preservation, and quality constraints, developed next.
@@
All-hex meshing of complex CAD models begins with a sound geometric decomposition. Downstream schemes, including unstructured approaches (for example, pave and sweep), structured mappings (for example, map), and block-structured layouts, then generate conforming, high-quality hexahedra. The central difficulty is not meshing a single sweepable or mappable block, but discovering a valid global decomposition and an executable sequence of operations that (i) respect sweepability constraints, (ii) preserve curvature and sharp features, and (iii) maintain element quality. Decades of practice have codified this workflow that provides robust predicates and operations, yet the pivotal decomposition step remains largely manual. This gap motivates AI methods that can capture expert strategies, reason over long action horizons, and adapt decisions to complex geometry.
@@
Recent work has begun to push in this direction. Patel et al. [Patel2021AutoHex] introduced Auto-Hex, which integrates geometric reasoning with RL. Starting from a boundary representation (B-Rep) and chordal-axis skeleton, the system identifies features such as T-junctions, sweep directions, and through-holes. An RL policy then proposes webcut actions conditioned on these cues, with feedback from the meshing environment guiding the agent to favor decompositions that yield hex or hex-dominant meshes. In related 2D work discussed earlier, DiPrete et al. [diprete2024reinforcement] used CAD shape testbeds to study RL-driven cut sequencing and reward design, while Zhang et al. [Zhang2025CADRLHex] frame 3D decomposition as a sequential decision problem and employ Soft Actor-Critic (SAC) to generate sweepable sub-volumes, showing promising decompositions on complex solids while highlighting open challenges in reward shaping and mesh quality.
@@
In parallel, supervised learning has been explored as another route to automation. Quadros et al. [Quadros2025Bipartite, QuadrosIMR2026] reduce a general 3D CAD model (outer BRep “skin”) to a compact 1D bipartite, heterogeneous graph derived from a 2D chordal axis skeleton (CAT), then apply a heterogeneous GNN to predict decomposition operations. Each node (surface or curve) carries geometric descriptors (e.g., area, length, centroid, thickness), topological cues (valency, interior loops), and meshing indicators (principal sweep directions), and the pipeline maintains bidirectional maps between BRep, CAT, and graph so predictions can be emitted as executable WebCut strings. Because labeled decompositions are costly, training is performed on randomized, synthetic CAD with multi-action labels; on held-out lattices the approach attains $98.72\%$ per action accuracy. Finally, they demonstrate effective performance on out-of-distribution geometries, achieving $85\%$ accuracy for WebCut predictions on complex industrial 3D CAD models.
Beyond aggregate accuracy, a key capability of this representation is that predictions can be traced back to local geometric and topological evidence in the graph. By propagating information over multi-hop neighborhoods, the GNN can learn recurring motifs that correlate with useful decomposition operations, such as cuts that expose sweep directions or isolate regions that are difficult to mesh. In practice, this shifts the workflow from manual exploration of many possible operations to a short list of plausible, high-impact candidates that can be applied directly through the mapped WebCut actions. Ongoing work extends this supervised framework to targeted CAD families and couples it with reinforcement learning, in the spirit of Auto-Hex agents, to learn action sequencing on general geometries without relying exclusively on labeled data.
Successful all-hex meshing is rarely a linear pipeline. Instead, it behaves like a tightly coupled loop in which decisions made early are repeatedly revisited as later checks expose new constraints. Practitioners typically cycle through physics-aware cleanup and defeaturing, specification of parting or symmetry, geometry decomposition via WebCuts, imprint and merge, sweepability checks and interval assignment, mesh generation, and quality assessment with localized repair, before attribution and model handoff. In practice, downstream checks frequently force returns to earlier stages. For example, imprint operations can introduce slender faces that require revised cuts, residual subvolumes may fail sweep checks and trigger alternative decompositions, and localized quality failures often motivate targeted changes to the cut plan rather than global retuning.
@@ 9 Volumetric Parameterizations
Volumetric parameterization maps a solid 3D domain to a simpler parametric volume, often a unit cube, a set of cuboids, or a polycube. In simulation workflows, this mapping is not an abstract convenience. It is a bridge from geometric shape to analysis-ready structure. For isogeometric analysis (IGA), many spline constructions rely on tensor-product bases, so a hexahedral control mesh and an analysis-suitable volumetric parameterization are closely linked [cottrell_isogeometric_2009]. For complex geometries, producing a bijective map with low distortion remains difficult, and that difficulty directly impacts numerical stability and accuracy in downstream analysis [zhang2016geometric].
Unlike block-structured meshing, which decomposes a geometry into discrete blocks to be meshed locally, volumetric parameterization seeks a continuous global map. When it succeeds, it enables two outcomes that are hard to obtain by purely local decisions: (i) all-hex meshes that act as control nets for spline construction, and (ii) tensor-product spline volumes whose structure aligns naturally with IGA discretizations [cottrell_isogeometric_2009]. When it fails, the failure modes are often decisive for simulation, including foldovers (negative Jacobians), excessive distortion, or singular structures that propagate into poor element quality and ill-conditioned systems.
A common theme in classical approaches is to enforce alignment to geometry and boundaries while also controlling injectivity and distortion. CubeCover [Nieser2011CubeCover] uses volumetric frame fields together with integer-grid maps to construct boundary-aligned parameterizations, illustrating how discrete constraints and continuous maps must be balanced to avoid inversions. Another widely used route is polycube-based parameterization, which deforms the input into a polycube domain and then pulls back a regular grid to obtain a structured hex layout [Gregson2011Polycube, Fang2016ClosedFormPolycube]. Recent work improves robustness and consistency through intrinsic formulations [Mandad2022IntrinsicPolycubes] and quantization strategies [Brueckler2022Quantization], aiming to better control alignment while reducing artifacts that degrade hex quality. Despite progress, these pipelines still face recurring bottlenecks on engineering models: layout decisions can be sensitive, optimization can be expensive, and user guidance is often needed to avoid poor decompositions and distorted regions [zhang2013challenges].
In IGA and geometry processing, parameterization is often framed as constructing spline-ready volumes rather than only extracting elements. Early approaches convert unstructured quad and hex meshes into T-spline surfaces and volumes [wenyan2010a, Wenyan2011a]. Harmonic and Laplacian-based formulations with Dirichlet boundary conditions have also been used to define mappings [Xia2010], supporting solid T-splines for genus-0 [zhang_solid_2012, Wenyan2013c] and higher-genus domains [LeiLiu2012a]. Polycube decomposition has been particularly influential for IGA because it provides a practical route to multi-cuboid parameter domains that can be parameterized by B-splines [LZH2014, Liu2015, HZ2016CMAME, HZL2017]. Alongside this, a range of spline technologies have been incorporated into analysis pipelines, including weighted T-splines [LeiLiu2015a], truncated T-splines [ref:wei16a], and truncated hierarchical B-splines [wei17a, wei2018blended]. These ideas have also reached commercial and production settings [Lai2016, Lai2017, yu2022hexgen], which underscores both the importance of the problem and the need for robust automation.
Across these families, the recurring challenges are consistent: computational cost, sensitivity to initialization and layout decisions, user dependence, and difficulty guaranteeing quality. For many geometries that arise in engineering, ensuring bijectivity while keeping distortion low is the central obstacle. When the mapping introduces poor elements, downstream optimization is often limited, and the resulting representations may be unsuitable for stable simulation.
AI is a good fit for volumetric parameterization for the same reason it is a good fit for earlier steps in the pipeline covered in this review. Practical parameterization depends on discrete, high-level decisions that are difficult to encode with fixed heuristics, including layout selection, surface labeling, and choosing boundary correspondences that balance alignment with distortion. Learning-based methods can propose these decisions quickly, while geometry-aware optimization can enforce smoothness and validity constraints.
DeepShape [DeepShapeCode] voxelizes the input and uses convolutional networks to establish correspondence with voxelized polycubes, reducing reliance on hand-tuned heuristics in the mapping process. Evocube [Dumery2022] explores evolutionary search to label faces and guide polycube construction, illustrating a complementary direction where AI explores a large discrete decision space that is otherwise difficult to navigate reliably. More recently, DL-Polycube [yu2025dlpolycube] predicts a polycube structure and uses it to guide surface segmentation, then combines octree subdivision, parametric mapping, and quality improvement to produce an all-hex mesh and an IGA-ready volumetric spline representation. This end-to-end pipeline is representative of how learning can reduce manual effort in polycube construction while keeping the downstream mapping and improvement stages grounded in classical geometry processing.
A second AI direction is to represent the parameterization using neural functions, which can serve as flexible approximators for complex mappings while incorporating geometric constraints. Penalty-based formulations remain important for controlling distortion and discouraging foldovers [Ji2022], and neural models can be used either to provide strong initializations or to learn components of the mapping that are expensive to derive from scratch. The integral parameterization approach in [Zhan2025] uses neural networks to represent an inverse parameterization while enforcing boundary correspondence and interior mapping, directly targeting the bijectivity and distortion tradeoffs that limit classical pipelines. Diffusion-based polycube generation has also been explored as a generative route to robust layouts, with DDPM-Polycube [yu2025DDPM] modeling polycube creation as a denoising process for producing high-quality hex meshes and volumetric splines.
From a readiness standpoint, the most reliable near-term pattern is hybrid: learning proposes layouts, segmentations, and initial correspondences quickly, then geometry-aware optimization enforces smoothness and validity. The key open gap for broad adoption in simulation is not only speed, but predictable behavior across diverse CAD topologies and feature scales [zhang2013challenges]. This places a premium on verification strategies for injectivity, robust handling of thin regions and complex topology, and benchmarking on engineering-relevant model collections. As these pieces mature, AI-empowered volumetric parameterization is positioned to reduce the cost of building analysis-suitable spline volumes and all-hex control meshes, expanding the practical reach of IGA in applications such as biomedicine, materials, and engineering design [cottrell_isogeometric_2009].
Contributed by Y. Zhang
@@ 10 Parallel Mesh Generation
As simulation fidelity and geometric complexity rise, large-scale FEA increasingly hinges on parallel mesh generation to sustain simulations with hundreds of millions of elements [chrisochoides2006]. Misconfigured runtimes can waste substantial wall time at this scale, so choosing the right parameters matters as much as the mesher itself. Both data and computation must be distributed over many processors in an HPC environment. Modern parallel meshers are indispensable, but they expose numerous runtime controls such as block size, work-partitioning strategy, NUMA affinity, and communication thresholds, and these choices have first-order effects on scalability and wall time. Selecting an optimal parameter vector $p$ is a nonlinear and architecture-sensitive problem that has traditionally depended on expert heuristics or costly trial and error.
@@
Analytic performance models exist, but they typically rest on simplifying assumptions and generalize poorly across architectures and problem families [barker2005practical]. Empirical tuning is more faithful but costly: each evaluation consumes a full HPC submission, and queue delays often dominate turnaround time. Collecting even a few dozen labeled runs can take weeks or even several months in production settings [angelopoulos2018]. These constraints impede routine optimization when rapid results are needed.
@@
Artificial intelligence offers a complementary path. In particular, supervised learning can train surrogate models that map parameter vectors to execution time [gamatie2019, ipek2005]. Training shifts cost up front. Once learned, the surrogate supports fast what-if exploration over unseen configurations without consuming additional HPC cycles, which enables wide synthetic searches of the parameter space.
As one illustrative case study, compact feedforward neural networks were trained as surrogate models to predict the execution time of a general-purpose parallel mesher from a parameter vector $p$ using small real-world datasets [garner2025scalability]. In that study, experiments were reported for two execution environments: a shared-memory cc-NUMA system with 172 runs [garner2025numa], and a distributed-memory setting with 68 runs [garner2025dist]. The distributed-memory results drew on representative node configurations including (i) dual-socket AMD EPYC 7763 nodes @ $2.45$ GHz ($64$ cores per socket, $128$ total slots) with $256$ GB memory, and Intel Xeon nodes such as (ii) dual-socket Xeon Gold 6148 @ $2.40$ GHz ($20$ cores per socket, $40$ total slots) with $384$ GB memory, and (iii) dual-socket Xeon E5-2698 v3 @ $2.30$ GHz ($16$ cores per socket, $32$ total slots) with $128$ GB memory. Reported results indicated that hierarchical load balancing was particularly beneficial on cc-NUMA platforms, consistent with the importance of locality for memory-bound stages [garner2025numa, garner2025dist].
In the same case study, the trained models scored 100,000100{,}000 synthetic configurations in under $0.5$ s per case, turning weeks of batch exploration into minutes. On held-out tests, the models reached a Mean Absolute Percentage Error (MAPE) of $2.13\%$ on the shared-memory system and $5.68\%$ on the distributed-memory system [tsolakis2024, garner2025dist]. This accuracy was sufficient to select settings that reduced total execution time. In practical terms, the primary bottleneck in HPC workflows is often cumulative queue wait time, and the study estimated an about $98\%$ reduction in total time-to-solution relative to a conservative manual search that would require tens or hundreds of queued jobs.
The core challenge in optimizing performance lies in the complexity of modern meshing codes. The workflow is decomposed into multiple, tightly coupled, parallel stages, where the optimal parameter choice for one stage is often dependent on the choices made for others [chrisochoides2006].
Domain Decomposition: The initial geometric domain is partitioned into smaller subdomains for concurrent processing by different processors [chrisochoides2006]. The decomposition strategy significantly influences communication patterns and load balancing, where a poor choice can cause substantial overhead [barker2005practical, thomadakis]. In addition to methods like recursive coordinate bisection, geometry-based techniques such as PQR (a sorting-based method utilizing a curvilinear coordinate system that binds to the boundary for partitioning elements) were also used in specific test cases [chrisochoides1994, garner2025dist]. The performance impact is determined by the surface-to-volume ratio of the subdomains.
Parallel Meshing and Refinement: Each subdomain is meshed simultaneously using core methods like Delaunay triangulation [Foteinos_JPDC_2014]. Adaptive isotropic mesh generation employs parameters such as $\delta$ (determines element size) [Foteinos_JPDC_2014] along with nthreadsnthreads and an adaptation heuristic such as adaptconstadapt-const [tsolakis2024, garner2025dist]. Parameters control criteria such as point insertion and topological modifications (e.g., local reconnection flips) [tsolakis2024].
Dynamic Load Balancing: A runtime system dynamically manages the distribution of work to mitigate computational load imbalance [thomadakis]. The implemented hierarchical load balancing model specifically targets cc-NUMA architectures. It prioritizes assigning workloads to threads that are pinned to CPU cores within the shortest NUMA node distance, effectively reducing costly remote memory accesses [garner2025numa, garner2025dist].
The surrogate model uses a feature vector $\mathbf{x}$ comprised of algorithmic, quality, and parallel parameters to predict the execution time. This parameter space is large and complex, as summarized in Table 2, drawing from the general problem structure.
Table 2: Examples of Key Parameter Categories Influencing Parallel Mesh Generation Performance
Category Examples of Parameters Primary Impact on Performance
Algorithmic Meshing technique (Delaunay) Controls efficiency and quality guarantees.
Execution model (eg. speculative [Foteinos_JPDC_2014]) Affects conflict resolution overhead (e.g., rollbacks).
Quality & Fidelity Element quality criteria Determines stopping criteria.
Sizing control method Dictates final mesh density and complexity.
Surface fidelity control [Drakopoulos2024] Deviation tolerance from input geometry.
Parallelism Domain decomposition method Influences communication patterns.
(Data/Task) Load balancing scheme Minimizes processor idle time.
Task granularity [Tsolakis2022] Ratio of computation to communication.
System/Runtime Inter-processor communication Time spent exchanging boundary data.
Load imbalance / quality Uneven work or poor initial partitioning.
Number/type of processors/cores [Thomadakis2022] Total parallel resource count.
@@
Beyond surrogates, recent work uses learned modules to dynamically decide about: (1) load balancing, where trained classifiers select domain-decomposition settings and reduce time-to-solution compared to static heuristics for production workflows like snappyHexMesh [Gangopadhyay2025ParCFD]; (2) decomposition and partitioning, where reinforcement learning (RL) with graph neural networks (GNNs) learns partitioners that compete with multilevel methods while providing a trainable interface to work distribution and communication footprint [Gatti2022RLGNN]; and (3) adaptation, where adaptive mesh refinement (AMR) is posed as a sequential decision problem, and learned policies make locality-aware refinement choices that scale across partitions and generalize across PDEs and mesh sizes [Yang2023RLAMR].